Solve each of the following systems by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
x+5y=10
-2x-10y=-20

Answer by stanbon(75887)

1st: x + 5y=10
2nd: -2x-10y=-20
----------------
Multiply 1st by -2 to get
-2x-10y = -20
------------------
The two equations are essentially the same.
The system is "dependent"
==============
Cheers,
Stan H.